how do u solve for x if the perimeter of the rectangle is x
EX. perimeter is 51 the sides we know are 8x-3 and 2x+11

Answer by unlockmath(1688)

Hello,
The formula for perimeter is P=2l+2w ---l is length and w is width. So just plug in what we have:
2(8x-3)+2(2x+11)=51
Expand out to:
16x-6+4x+22=51
Combine like terms:
20x+16=51
Subtract 16 and divide 20 into both sides to get:
x=35/20 or 1 3/4
Make sense?
RJ
www.math-unlock.com